# Minimax and adaptive estimation of general linear functionals under sparsity

### v1

@@ C.2 Additional proof of the lower bounds in  Theorem 1
$1+\chi^{2}(P_{2}\|P_{1})$ $=\int_{\mathbb{R}^{d}}\Bigg[\sum_{j=1}^{d}\pi_{j}\frac{f_{\alpha}^{(0)}\!\left(x_{j}-c\frac{\lambda_{o}}{\eta_{j}}\right)}{f_{\alpha}^{(0)}(x_{j})}\Bigg]^{2}\,dP_{1}$
$=\sum_{i\neq j}\pi_{i}\pi_{j}\int_{\mathbb{R}^{2}}f_{\alpha}^{(0)}\!\left(x_{i}-c\frac{\lambda_{o}}{\eta_{i}}\right)f_{\alpha}^{(0)}\!\left(x_{j}-c\frac{\lambda_{o}}{\eta_{j}}\right)\,dx_{i}\,dx_{j}$
$\quad+\sum_{j=1}^{d}\pi_{j}^{2}\int_{\mathbb{R}}\frac{f_{\alpha}^{(0)}\!\left(x_{j}-c\frac{\lambda_{o}}{\eta_{j}}\right)^{2}}{f_{\alpha}^{(0)}(x_{j})}\,dx_{j}$
@@
C.3 Proof of the lower bound in Theorem 8
@@
C.4 Proofs of lemmas in Section 8.4
@@
where $h_{j}^{2}=\int\left(\sqrt{f_{0}(x)}-\sqrt{\frac{f_{0}((x+\pi_{j}\gamma_{j})/\sigma_{j})}{\sigma_{j}}}\right)^{2}dx$. Let $I_{1}(t)$ and $I_{2}(t;\sigma_{j})$ be the Fisher information of the scale family $\{f_{0}(x/t)/t:t>0\}$ and of the location family $\{f_{0}((x+t)/\sigma_{j})/\sigma_{j}:t\in[0,1]\}$, respectively. We have
@@
$\frac{1}{2}h_{j}^{2}$ $\leq\int\left(\sqrt{f_{0}(x)}-\sqrt{\frac{f_{0}(x/\sigma_{j})}{\sigma_{j}}}\right)^{2}dx+\int\left(\sqrt{\frac{f_{0}(x/\sigma_{j})}{\sigma_{j}}}-\sqrt{\frac{f_{0}((x+\pi_{j}\gamma_{j})/\sigma_{j})}{\sigma_{j}}}\right)^{2}dx$
@@ D.1 Homogeneous loading vector
    $\sum_{j\leq(s^{2}\wedge d)}\eta_{j}^{2}\;=\;s^{2}\wedge d.$

### v2

@@ C.3 Additional proof of the lower bounds in  Theorem 1
$1+\chi^{2}(P_{2}\|P_{1})$ $=\int_{\mathbb{R}^{d}}\Bigg[\sum_{j=1}^{d}\pi_{j}\frac{f_{\alpha}^{(0)}\!\left(x_{j}-c\frac{\lambda_{o}}{\eta_{j}}\right)}{f_{\alpha}^{(0)}(x_{j})}\Bigg]^{2}\,{\rm d\,}P_{1}$
$=\sum_{i\neq j}\pi_{i}\pi_{j}\int_{\mathbb{R}^{2}}f_{\alpha}^{(0)}\!\left(x_{i}-c\frac{\lambda_{o}}{\eta_{i}}\right)f_{\alpha}^{(0)}\!\left(x_{j}-c\frac{\lambda_{o}}{\eta_{j}}\right)\,{\rm d\,}x_{i}\,{\rm d\,}x_{j}$
$\quad+\sum_{j=1}^{d}\pi_{j}^{2}\int_{\mathbb{R}}\frac{f_{\alpha}^{(0)}\!\left(x_{j}-c\frac{\lambda_{o}}{\eta_{j}}\right)^{2}}{f_{\alpha}^{(0)}(x_{j})}\,{\rm d\,}x_{j}$
@@
C.4 Proof of the lower bound in Theorem 8
@@
C.5 Proofs of lemmas in Section C.1
@@
where $h_{j}^{2}=\int\left(\sqrt{f_{0}(x)}-\sqrt{\frac{f_{0}((x+\pi_{j}\gamma_{j})/\sigma_{j})}{\sigma_{j}}}\right)^{2}{\rm d\,}x$. Let $I_{1}(t)$ and $I_{2}(t;\sigma_{j})$ be the Fisher information of the scale family $\{f_{0}(x/t)/t:t>0\}$ and of the location family $\{f_{0}((x+t)/\sigma_{j})/\sigma_{j}:t\in[0,1]\}$, respectively. We have
@@
$\frac{1}{2}h_{j}^{2}$ $\leq\int\left(\sqrt{f_{0}(x)}-\sqrt{\frac{f_{0}(x/\sigma_{j})}{\sigma_{j}}}\right)^{2}{\rm d\,}x+\int\left(\sqrt{\frac{f_{0}(x/\sigma_{j})}{\sigma_{j}}}-\sqrt{\frac{f_{0}((x+\pi_{j}\gamma_{j})/\sigma_{j})}{\sigma_{j}}}\right)^{2}{\rm d\,}x$
@@ D.1 Homogeneous loading vector
    $\sum_{j\leq(s^{2}\wedge d)}\eta_{j}^{2}\;=\;j_{3}(s)=\big\lceil s^{2}\log^{2/\alpha}(ed/s)\big\rceil\wedge d.$